# UniGAP: A Universal and Adaptive Graph Upsampling Approach to Mitigate Over-Smoothing in Node Classification Tasks

### v1

@@ Abstract.
In the graph domain, deep graph networks based on Message Passing Neural Networks (MPNNs) or Graph Transformers often cause over-smoothing of node features, limiting their expressive capacity. Many upsampling techniques involving node and edge manipulation have been proposed to mitigate this issue. However, these methods often require extensive manual labor, resulting in suboptimal performance and lacking a universal integration strategy. In this study, we introduce UniGAP, a universal and adaptive graph upsampling technique for graph data. It provides a universal framework for graph upsampling, encompassing most current methods as variants. Moreover, UniGAP serves as a plug-in component that can be seamlessly and adaptively integrated with existing GNNs to enhance performance and mitigate the over-smoothing problem. Through extensive experiments, UniGAP demonstrates significant improvements over heuristic data augmentation methods across various datasets and metrics. We analyze how graph structure evolves with UniGAP, identifying key bottlenecks where over-smoothing occurs, and providing insights into how UniGAP addresses this issue. Lastly, we show the potential of combining UniGAP with large language models (LLMs) to further improve downstream performance. Our code is available at: https://github.com/wangxiaotang0906/UniGAP
@@ 1. Introduction
Graph data is ubiquitous in the real world applications like recommender systems (van den Berg et al., 2017; Ying et al., 2018), traffic analysis (Zheng et al., 2020; Guo et al., 2019), and drug discovery (Gaudelet et al., 2021). Graph neural networks (GNNs) (Defferrard et al., 2016; Velickovic et al., 2018; Hamilton et al., 2017) are powerful tools for handling graph-structured data, demonstrating significant success in recent years. In contrast to computer vision and natural language processing, where deep models often lead to significant improvements (He et al., 2016), in the graph domain, increasing the number of model layers can cause node features to converge exponentially towards the same vector, a phenomenon known as “over-smoothing” (Rusch et al., 2023; Keriven, 2022; Chen et al., 2020a). This convergence diminishes the model’s expressive power and adversely affects its performance.
@@
Previous researches have explored a variety of strategies to mitigate the over-smoothing problem, including manipulating graph topology (Chen et al., 2020a; Rong et al., 2020; You et al., 2020), refining model structure (Kipf and Welling, 2017; Xu et al., 2018), and employing dynamically learning (Zhang et al., 2023; Zou et al., 2023). Recently, graph upsampling, a lightweight graph topology manipulation strategy, has proven effective in complementing and enhancing other techniques. For instance, AdaEdge (Chen et al., 2020a) heavily relys on “homophily” assumption and alleviate over-smoothing by increasing intra-class and decreasing inter-class connections based on the model’s classification outcomes. HalfHop (Azabou et al., 2023) proposes a heuristic method of randomly adding nodes to slow down message propagation. However, these approaches rely on heuristic strategies that require prior knowledge and significant human intervention, often resulting in a sub-optimal performance and lack of interpretability, which makes it challenging to understand their effectiveness or rationale. Moreover, there is currently no universal framework that describes and encompasses these methods, despite their common manipulation of graph topology and interconnected effects.
@@
To address the aforementioned challenges, we introduce UniGAP, a universal and adaptive graph upsampling methodology. Specifically, it offers a universal framework that subsumes various graph upsampling techniques as its variants. For instance, methods such as HalfHop, AdaEdge, and other upsampling techniques (Wu et al., 2021; Liu et al., 2022) can be considered specific variants of UniGAP (Section 3.4). Furthermore, different from existing methods that employ a heuristic upsampling strategy, UniGAP adaptively calculates the probability of inserting an intermediate node for each pair of adjacent nodes in the original graph, generating a new upsampled graph through a differentiable sampling technique. Subsequently, UniGAP iteratively refines these sampling probabilities based on evaluation metrics (e.g., accuracy or smoothness) of the upsampled graph in downstream tasks. This iterative process strives to yield an optimally upsampled graph suitable for downstream applications. Moreover, UniGAP enhances the interpretability of the over-smoothing issue within graph structure learning. By examining the augmented graph that achieves optimal performance, we can precisely identify where intermediate nodes are inserted. This analysis provides valuable insights into the root causes of the over-smoothing problem. Lastly, we demonstrate the potential of integrating LLMs with UniGAP to further improve downstream performance.
@@
  • We propose UniGAP, a universal and adaptive graph upsampling approach to mitigate over-smoothing in node classification and enhance downstream performance as a plug-in. Additionally, UniGAP can encompass various graph upsampling techniques as its variants.
@@ 2.1. Over-Smoothing Problem in GNNs
Over-smoothing represents a prevalent research issue within the realm of graphs, garnering significant attention recently. This phenomenon substantially impedes the expressiveness of GNNs as the number of propagation layers increases, resulting in node features becoming progressively more indistinguishable. To assess the extent of over-smoothing, a variety of quantitative metrics have been established, including Mean Average Distance (MAD) (Chen et al., 2020b), Euclidean Distance (Zhou et al., 2020a), Dirichlet Energy (Rusch et al., 2023), and others.
Previous research (Cai and Wang, 2020; Zhang et al., 2024; Rusch et al., 2023) has explored various strategies to mitigate the over-smoothing problem. These solutions can be broadly categorized into three main approaches: manipulating graph topology, refining model structure, and dynamic learning. Firstly, manipulating graph topology involves modifying the original graph structure using dropout techniques (e.g., DropEdge (Rong et al., 2020), DropNode (Feng et al., 2021), DropMessage (Fang et al., 2023)), graph upsampling (Azabou et al., 2023; Chen et al., 2020a), or others (Han et al., 2022). These methods aim to reduce message passing or enhance graph structure to alleviate over-smoothing. Secondly, refining model structure includes modifying message passing mechanisms (Chien et al., 2021) or utilizing some model tricks like residual connections (Chen et al., 2020c), multi-hop information (Xu et al., 2018) to mitigate over-smoothing. Lastly, dynamic learning strategies dynamically adjusting the learning process of the model. For instance, DAGNN (Zhou et al., 2020b) calibrates a retention score to preserve representations in each layer, DGSLN (Zou et al., 2023) propose a graph generation module to optimize the graph structure, and ADGCN (Zhang et al., 2023) employs attention for selective cross-layer information aggregation.
@@ 2.2. Graph Augmentations
Graphs, as non-Euclidean data, primarily consist of entities and the relations between them. Therefore, augmentation techniques for graph data are inherently more complex compared to other domains such as computer vision and natural language processing. Graph augmentations aim to enhance model generalization, improve performance, and mitigate over-smoothing. These augmentations can be categorized into three main types: node-level, edge-level, and graph-level augmentations. Node-level augmentations focus on individual nodes. For example, DropNode (You et al., 2020) randomly sets some nodes as zero vectors, FeatureMask (Zhu et al., 2020a) masks partial features of each node randomly, and FLAG (Kong et al., 2022) adds a learnable perturbation to each node. Edge-level augmentations manipulate edges within the graph. For instance, DropEdge (Rong et al., 2020) and EdgeMask (You et al., 2020) randomly remove existing edges based on specified probabilities. AdaEdge (Chen et al., 2020a) relies on a heuristic approach by adjusting intra-class and inter-class connections based on model classification outcomes. GAug (Zhao et al., 2021) uses neural edge predictors to estimate edge probabilities. Graph-level augmentations modify both nodes and edges simultaneously. GraphCL (You et al., 2020) combines various augmentation techniques to modify the original graph. Additionally, GCA (Zhu et al., 2021) proposes adaptive augmentation approaches based on metrics such as the connectivity of target nodes. HalfHop (Azabou et al., 2023) suggests adding nodes to existing edges randomly to slow down message propagation and mitigate over-smoothing.
@@ 3. UniGAP: A Universal and Adaptive Graph Upsampling Method
In this section, we will start with the notations we use throughout the rest of the paper in Section 3.1. Subsequently, we will provide an overview of our method in Section 3.2. Then we will introduce each component in our method for details in Section 3.3. Finally, we show UniGAP as a universal framework in Section 3.4 and discuss UniGAP from a theoretical perspective in Section 3.5.
Figure 1. Framework of UniGAP. There are four main components in UniGAP, including Trajectory Precomputation, Multi-View Condensation (MVC) Encoder, Adaptive Upsampler, and the downstream model. Trajectory Precomputation needs to be performed only once in each task, then MVC, Adaptive Upsampler, and downstream are co-trained to obtain the optimal graph structure.
3.1. Notations
@@
In this work, our focus is on node classification, which involves learning a graph predictor $g_{\theta}:\mathcal{G}\mapsto\mathcal{Y}$ that maps instance $G=(\mathcal{V},\mathcal{E},{A},{X},{Y})\in\mathcal{G}$ to label ${Y}\in\mathcal{Y}$, where $\mathcal{V}$ and $\mathcal{E}$ denote the sets of nodes and edges, respectively. ${X}\in\mathbb{R}^{N\times d}$ represents the node attribute feature matrix with $N$ nodes and $d$ attributes. ${A}\in\mathbb{R}^{N\times N}$ is the adjacency matrix. Additionally, we denote the diagonal degree matrix as $D\in\mathbb{R}^{N\times N}$, and the matrices with added self-loops as $\tilde{A}\in\mathbb{R}^{N\times N}$ and $\tilde{D}\in\mathbb{R}^{N\times N}$, respectively. The graph Laplacian with self-loops is defined as $\tilde{L}=\tilde{D}-\tilde{A}$. For normalization, the symmetrically normalized adjacency matrix is $\hat{A}=\tilde{D}^{-\frac{1}{2}}\tilde{A}\tilde{D}^{-\frac{1}{2}}$, and the symmetrically normalized graph Laplacian is $\hat{L}=\tilde{D}^{-\frac{1}{2}}\tilde{L}\tilde{D}^{-\frac{1}{2}}$. For clarity in notations, we will use $A$ and $L$ to denote the original graph, regardless of whether it is normalized or unnormalized. The augmented graph, which will be introduced later, will be represented by $\hat{A}$ and $\hat{L}$.
3.2. Overview of UniGAP
In this section, we will provide an overview of our pipeline, with a detailed introduction of our method available in subsequent sections. UniGAP consists of four main components: Trajectory Computation, Multi-View Condensation (MVC) Encoder, Adaptive Upsampler, and the downstream model.
Initially, we compute trajectories (multi-hop information) of the input graph using our Trajectory Precomputation module, denoted as $T=\mathcal{P}(G)$, where $T\in\mathbb{R}^{L\times N\times d}$ represents the trajectories and $\mathcal{P}(\cdot)$ denotes the frozen trajectory precomputation module. These trajectories $T$ are then processed by a tunable Multi-View Condensation (MVC) Encoder $\mathcal{F}_{\omega}(\cdot):\mathbb{R}^{L\times N\times d}\mapsto\mathbb{R}^{N\times d}$ to obtain condensed information, denoted as $\hat{T}=\mathcal{F}_{\omega}(T)$. Next, the condensed information $\hat{T}$ is used to compute the node insertion probabilities between each pair, denoted as $P=\mathcal{U}_{\phi}(\hat{T})$, where $\mathcal{U}_{\phi}:\mathbb{R}^{N\times N}\mapsto\mathbb{R}^{E\times 1}$ represents adaptive upsampler module. Building upon this, a differentiable upsampling operation $\mathcal{S}$ is employed to obtain the augmented graph $\hat{G}=(\hat{A},\hat{X})=\mathcal{S}(P)$. Finally, the augmented graph will be input into the downstream model $\mathcal{D}_{\psi}$ for specific tasks. Using node classification as an example, the cross-entropy loss will be computed to iteratively optimize the previous learnable modules. In subsequent epochs, the refined downstream model will compute trajectories, with other steps remaining consistent throughout the iteration. The pipeline of our method is illustrated in Figure 1 and Algorithm 1. Next, we will delve into each component in detail.
3.3. UniGAP as An Adaptive Upsampling Framework
3.3.1. Trajectory Computation
Trajectories encompass multi-view information of each node, including one-hop, two-hop, and higher-order information, which are essential for subsequent modules. As the output of node features at each layer of the model, they implicitly reflect the gradual process of over-smoothing as the number of layers increases. Our aim is to derive a high-dimensional representation of the over-smoothing process from the trajectories, thereby enabling us to devise strategies to mitigate over-smoothing. In this section, we will elucidate the process of trajectory computation.
@@ 3.3.1. Trajectory Computation
Let the trajectories be denoted as $T\in\mathbb{R}^{L\times N\times d}$, where $L$ represents the number of layers, $N$ denotes the number of nodes, and $d$ signifies the dimensionality of features. Trajectories include the concatenation of multi-view features (hidden states from each layer):
(1) $T=\operatorname{concat}({T}^{(1)},{T}^{(2)},\cdots,{T}^{(L)})\in\mathbb{R}^{L\times N\times d}$
where $\operatorname{concat}$ denotes channel-wise concatenation. $L$ is the number of model layers. $T^{(1)},\cdots,T^{(L)}$ represent multi-view features (1-hop, $\cdots$, L-hop information), we collect them through gathering the hidden representations in each layer:
(2) ${T}^{(l)}=\begin{cases}\operatorname{Norm}({H}^{(l)}),&\text{if}\ \operatorname{mod}(l,w)=0\\ {H}^{(l)},&\text{o.w.}\\ \end{cases}$
where the representation of each hop is denoted as $H^{(l)}=f^{(l)}(H^{(l-1)})\in\mathbb{R}^{N\times d}$, $f^{(l)}(\cdot)$ means $l$-th layer in trajectory computation model, and $\operatorname{Norm}({x}_{i})=\frac{{x}_{i}}{\left|\left|{x}_{i}\right|\right|_{2}}$ signifies a channel-wise $L2$ normalization, which aids in regulating the magnitude of features that may escalate exponentially with the number of iterations. $w$ is a hyper-parameter that determines the frequency of normalizing features.
@@
2) Non-parametric Message Passing. After removing the non-linear operator in graph neural networks, we can decouple the message passing ($AX$) and feature transformation ($XW$). However, in the initial stage, we have no access to the transformation weights $W$ unless pre-training methods are employed. Therefore, we adopt non-parameteric message passing to collect multi-view information, i.e., $AX,A^{2}X,\cdots A^{L}X$. This approach avoids nonlinear transformations, preserving maximum original feature, graph topology, and trajectories. The formulation is as follows:
@@
3.3.2. Multi-View Condensation
After obtaining trajectories (multi-view information), we aim to adaptively condense this multi-view information into compact features that reflect the node’s propensity to over-smoothing, while considering efficiency and adaptivity. To achieve this goal, we design a multi-view condensation (MVC) encoder that can adaptively condense trajectories $T\in\mathbb{R}^{L\times N\times d}$ into condensed trajectory features $\hat{T}\in\mathbb{R}^{N\times d}$. This reduces computational cost and adaptively adjusts the weights of information at each hop. We instantiate the MVC encoder using two strategies in this work: Trajectory-MLP-Mixer (TMM) and Trajectory-Transformer.

### v2

@@ Abstract.
In the graph domain, deep graph networks based on Message Passing Neural Networks (MPNNs) or Graph Transformers often cause over-smoothing of node features, limiting their expressive capacity. Many upsampling techniques involving node and edge manipulation have been proposed to mitigate this issue. However, these methods are often heuristic, resulting in extensive manual labor and suboptimal performance and lacking a universal integration strategy. In this study, we introduce UniGAP, a universal and adaptive graph upsampling framework to mitigate over-smoothing in node classification tasks. Specifically, we design an adaptive graph upsampler based on condensed trajectory features, serving as a plug-in component for existing GNNs to mitigate the over-smoothing problem and enhance performance. Moreover, UniGAP serves as a representation-based and fully differentiable framework to inspire further exploration of graph upsampling methods. Through extensive experiments, UniGAP demonstrates significant improvements over heuristic data augmentation methods in various datasets and metrics. We analyze how graph structure evolves with UniGAP, identifying key bottlenecks where over-smoothing occurs, and providing insights into how UniGAP addresses this issue. Lastly, we show the potential of combining UniGAP with large language models (LLMs) to further improve downstream performance. Our code is available at: https://anonymous.4open.science/r/UniGAP.
@@ 1. Introduction
Graph neural networks (GNNs) (Defferrard et al., 2016; Velickovic et al., 2018; Hamilton et al., 2017) are powerful tools for handling graph-structured data, demonstrating significant success in recent years (van den Berg et al., 2017; Ying et al., 2018; Zheng et al., 2020; Guo et al., 2019; Gaudelet et al., 2021). In contrast to computer vision and natural language processing, where deep models often lead to significant improvements (He et al., 2016), in the graph domain, increasing the number of model layers can cause node features to converge exponentially towards the same vector, a phenomenon known as “over-smoothing” (Rusch et al., 2023; Keriven, 2022; Chen et al., 2020a). This convergence diminishes the model’s expressive power and adversely affects its performance.
@@
Previous research has explored a variety of strategies to mitigate the over-smoothing problem, including manipulating graph topology (Chen et al., 2020a; Rong et al., 2020; You et al., 2020), refining model structure (Kipf and Welling, 2017; Xu et al., 2018), and employing dynamical learning (Zhang et al., 2023; Zou et al., 2023). Recently, graph upsampling, a lightweight graph topology manipulation strategy, has proven effective in complementing and enhancing other techniques. AdaEdge (Chen et al., 2020a) relies on prior knowledge encoded in a heuristic adaptation rule—namely, increasing intra-class connectivity and decreasing inter-class connectivity—while HalfHop (Azabou et al., 2023) depends on a predefined random insertion strategy. Such heuristic designs result in suboptimal performance and lack of interpretability, making it difficult to understand their effectiveness or rationale.
@@
To address the aforementioned challenges, we introduce UniGAP, a universal and adaptive graph upsampling methodology. Specifically, different from existing methods that employ a heuristic upsampling strategy, UniGAP adaptively calculates the probability of inserting an intermediate node for each pair of adjacent nodes in the original graph, generating a new upsampled graph through a differentiable sampling technique. Subsequently, UniGAP iteratively refines these sampling probabilities based on evaluation metrics (e.g., accuracy or smoothness) of the upsampled graph in downstream tasks. This iterative process strives to yield an optimally upsampled graph suitable for downstream applications. Additionally, UniGAP serves as a fully differentiable representation-based framework to inspire further exploration of graph upsampling methods, in which its components are all modular and can be replaced in different ways. Furthermore, UniGAP enhances the interpretability of the over-smoothing issue within graph structure learning. By examining the augmented graph that achieves optimal performance, we can precisely identify where intermediate nodes are inserted. This analysis provides valuable insights into the root causes of the over-smoothing problem. Lastly, we demonstrate the potential of integrating LLMs with UniGAP to further improve downstream performance.
@@
  • We propose UniGAP, a universal and adaptive graph upsampling approach to mitigate over-smoothing in node classification and enhance downstream performance as a plug-in. Additionally, UniGAP serves as a fully differentiable trajectory-based framework to inspire further exploration of graph upsampling methods.
@@ 2.1. Over-Smoothing Problem in GNNs
Over-smoothing represents a prevalent research issue within the realm of graphs, garnering significant attention recently. This phenomenon substantially impedes the expressiveness of GNNs as the number of propagation layers increases, resulting in node features becoming progressively more indistinguishable. To assess the extent of over-smoothing, a variety of quantitative metrics have been established, including Mean Average Distance (MAD) (Chen et al., 2020b), Euclidean Distance (Zhou et al., 2020b), Dirichlet Energy (Rusch et al., 2023), and others.
Previous research (Cai and Wang, 2020; Zhang et al., 2024; Rusch et al., 2023) has explored various strategies to mitigate the over-smoothing problem. These solutions can be broadly categorized into three main approaches: manipulating graph topology, refining the model structure, and dynamic learning. Firstly, manipulating graph topology involves modifying the original graph structure using dropout techniques (e.g., DropEdge (Rong et al., 2020), DropNode (Feng et al., 2021), DropMessage (Fang et al., 2023)), graph upsampling (Azabou et al., 2023; Chen et al., 2020a), or others (Han et al., 2022). These methods aim to reduce message passing or enhance graph structure to alleviate over-smoothing. Secondly, refining the model structure includes modifying message passing mechanisms (Chien et al., 2021) or utilizing some model tricks like residual connections (Chen et al., 2020c) and multi-hop information (Xu et al., 2018) to mitigate over-smoothing. Lastly, dynamic learning strategies dynamically adjust the learning process of the model. For instance, DAGNN (Zhou et al., 2020a) calibrates a retention score to preserve representations in each layer, DGSLN (Zou et al., 2023) proposes a graph generation module to optimize the graph structure, and ADGCN (Zhang et al., 2023) employs attention for selective cross-layer information aggregation.
@@ 2.2. Graph Augmentations
Graphs, as non-Euclidean data, primarily consist of entities and the relations between them. Therefore, augmentation techniques for graph data are inherently more complex compared to other domains such as computer vision and natural language processing. Graph augmentations aim to enhance model generalization, improve performance, and mitigate over-smoothing. These augmentations can be categorized into three main types: node-level, edge-level, and graph-level augmentations. Node-level augmentations focus on individual nodes. For example, DropNode (You et al., 2020) randomly sets some nodes as zero vectors, FeatureMask (Zhu et al., 2020b) randomly masks partial features of each node, and FLAG (Kong et al., 2022) adds a learnable perturbation to each node. Edge-level augmentations manipulate edges within the graph. For instance, DropEdge (Rong et al., 2020) and EdgeMask (You et al., 2020) randomly remove existing edges based on specified probabilities. AdaEdge (Chen et al., 2020a) relies on a heuristic approach by adjusting intra-class and inter-class connections based on model classification outcomes. GAug (Zhao et al., 2021) uses neural edge predictors to estimate edge probabilities. Graph-level augmentations modify both nodes and edges simultaneously. GraphCL (You et al., 2020) combines various augmentation techniques to modify the original graph. UGA (Zhuo et al., 2024) proposes a unified graph augmentation framework for generalized contrastive learning that automatically selects and composes suitable augmentations for different graphs and tasks. Additionally, GCA (Zhu et al., 2021) proposes adaptive augmentation approaches based on metrics such as the connectivity of target nodes. HalfHop (Azabou et al., 2023) suggests adding nodes to existing edges randomly to slow down message propagation and mitigate over-smoothing.
@@
2.3. Graph Structure Learning
Graph structure learning (GSL) has emerged as a complementary line of research to graph augmentation. While graph augmentations typically apply stochastic or rule-based transformations to generate alternative graph views for regularization or contrastive learning, GSL aims to explicitly learn an optimized graph structure jointly with node representations in an end-to-end manner. A large body of GSL methods focuses on learning or refining the adjacency matrix through edge rewiring, i.e., adding, deleting, or reweighting edges between existing nodes. Some approaches learn a latent graph from node features and optimize it together with GNN parameters for downstream prediction (Chen et al., 2020d; Franceschi et al., 2020; Shen et al., 2024). In parallel, a series of rewiring methods modifies the graph topology to alleviate over-smoothing or over-squashing, for example, by inserting long-range edges, sparsifying redundant connections, or reconfiguring edges based on curvature or edge-noise estimates (Dai et al., 2022; Arnaiz-Rodriguez et al., 2022). These methods typically operate in the space of edge-level operations and keep the set of nodes fixed.
Compared with these rewiring-based GSL approaches, UniGAP takes a different perspective by performing graph upsampling through node insertion on existing edges. Instead of only deciding whether an edge should be added, removed, or reweighted, UniGAP adaptively inserts intermediate nodes to effectively stretch selected edges, thus slowing down message passing along them.
@@ 3. UniGAP: A Universal and Adaptive Graph Upsampling Method
In this section, we introduce our proposed universal and adaptive graph upsampling method, i.e., UniGAP. Firstly, it computes the input graph’s trajectories (multi-hop information) using our Trajectory Precomputation module (Section 3.1). These trajectories are then processed by a tunable Multi-View Condensation (MVC) Encoder to extract condensed information (Section 3.2). Next, the condensed features are used to compute node insertion probabilities via the adaptive upsampler module (Section 3.3). Finally, the augmented graph is fed into the downstream model for specific tasks (Section 3.4). The pipeline and algorithmic description of UniGAP are shown in Figure 1 and Algorithm 1, respectively.
Figure 1. Overview of the UniGAP framework. (1) Given an input graph, the Trajectory Precomputation module first computes layer-wise node trajectories that capture how node representations evolve with depth. (2) These trajectories are fed into the Multi-View Condensation (MVC) Encoder, which aggregates multi-hop information into a compact, per-node representation. (3) The Adaptive Upsampler then uses these condensed features to predict where to insert intermediate nodes on edges, and a differentiable sampling step produces an augmented graph with updated topology and node features. (4) Finally, the augmented graph is passed to a downstream GNN for tasks such as node classification, and the task loss is backpropagated to jointly update the MVC encoder, the upsampler, and the downstream model. After a warm-up epoch, the refined downstream GNN is reused to recompute trajectories, and the process is iterated until convergence.
@@
Notations. In this work, our focus is on node classification, which involves learning a graph predictor $g_{\theta}:\mathcal{G}\mapsto\mathcal{Y}$ that maps instance $G=(\mathcal{V},\mathcal{E},{A},{X},{Y})\in\mathcal{G}$ to label ${Y}\in\mathcal{Y}$, where $\mathcal{V}$ and $\mathcal{E}$ denote the sets of nodes and edges, respectively. ${X}\in\mathbb{R}^{N\times d}$ represents the node attribute feature matrix with $N$ nodes and $d$ attributes. ${A}\in\mathbb{R}^{N\times N}$ is the adjacency matrix. Additionally, we denote the diagonal degree matrix as $D\in\mathbb{R}^{N\times N}$, and the matrices with added self-loops as $\tilde{A}\in\mathbb{R}^{N\times N}$ and $\tilde{D}\in\mathbb{R}^{N\times N}$, respectively. The graph Laplacian with self-loops is defined as $\tilde{L}=\tilde{D}-\tilde{A}$. For normalization, the symmetrically normalized adjacency matrix is $\hat{A}=\tilde{D}^{-\frac{1}{2}}\tilde{A}\tilde{D}^{-\frac{1}{2}}$, and the symmetrically normalized graph Laplacian is $\hat{L}=\tilde{D}^{-\frac{1}{2}}\tilde{L}\tilde{D}^{-\frac{1}{2}}$. For clarity in notations, we will use $A$ and $L$ to denote the original graph, regardless of whether it is normalized or unnormalized. The augmented graph, which will be introduced later, will be represented by $\hat{A}$ and $\hat{L}$.
3.1. Trajectory Computation
Considering the output of node features at each layer of the model ${H}^{(l)}$ implicitly reflects the gradual process of over-smoothing as the number of layers increases. We interpret these features as trajectories that model this process. Concretely, for each node, the sequence of its representations across layers forms a trajectory that encodes how it is affected by over-smoothing as depth grows. Building on this intuitive assumption, our framework leverages these trajectories to diagnose where and when over-smoothing occurs and to design strategies that adaptively modify the graph structure to mitigate its effects. Specifically, trajectory $T^{(l)}\in\mathbb{R}^{N\times d}$ refers to the hop information of each node, such as one-hop $H^{(1)}$, two-hop $H^{(2)}$, and higher-order features:
[…]
@@ 3.1. Trajectory Computation
These trajectories are essential for the functionality of subsequent modules, which will be optimized iteratively. Let the trajectories be denoted as $T=[{T}^{(1)},{T}^{(2)},\cdots,{T}^{(L)}]\in\mathbb{R}^{L\times N\times d}$, where $L$ represents the number of layers, $N$ denotes the number of nodes, and $d$ signifies the dimensionality of the features.
@@
2) Non-parametric Message Passing. After removing the non-linear operator in graph neural networks, we can decouple the message passing ($AX$) and feature transformation ($XW$). However, in the initial stage, we have no access to the transformation weights $W$ unless pre-training methods are employed. Therefore, we adopt non-parameteric message passing to collect multi-view information, i.e., $AX,A^{2}X,\cdots A^{L}X$. This approach avoids nonlinear transformations, preserving maximum original feature and pure graph topology information. The formulation is as follows:
@@
In Section 5.5, we experimentally explore various precomputation strategies to address the cold start problem in trajectory computation. Based on performance, we select Pretrained GNN as the precomputation approach for this work.
3.2. Multi-View Condensation
After obtaining trajectories (multi-view information), we aim to adaptively condense this multi-view information into compact features that reflect the node’s propensity to over-smoothing, while considering efficiency and adaptivity. To achieve this goal, we design a tunable multi-view condensation (MVC) encoder $\mathcal{F}_{\omega}(\cdot):\mathbb{R}^{L\times N\times d}\mapsto\mathbb{R}^{N\times d}$ that can adaptively condense trajectories $T\in\mathbb{R}^{L\times N\times d}$ into condensed trajectory features $\hat{T}\in\mathbb{R}^{N\times d}$. This reduces computational cost and adaptively adjusts the weights of information at each hop. In this work, we instantiate the MVC encoder using two strategies: Trajectory-MLP-Mixer (TMM) and Trajectory-Transformer (TT).